# Large Language Models Are Human-Like Internally

### v1

@@ 3.1 Human Data
To minimize confounding factors between stimuli data and human measures, we included datasets with multi-layered annotations across multiple human measures. These include the Natural Stories Corpus Futrell et al. (2021) with SPR and MAZE data Boyce and Levy (2023), ZuCO corpus Hollenstein et al. (2018) with FPGD and N400 data, UCL Corpus Frank et al. (2013) annotated with SPR, FPGD, and N400 data Frank et al. (2015), and filler sentences from Hahn et al. (2022) annotated with SPR, FPGD, and MAZE data Vasishth et al. (2010); Hahn et al. (2022). In particular, the FPGD and N400 data in ZuCO were simultaneously recorded from the same human subjects, which likely minimized confounding factors.
Figure 2: Relationships between layer depth (x-axis) and $\Delta$LL (y-axis) for each LM in two datasets: FPGD in DC and SPR in NS. The graphs are separated by model families and data. The best/last layer is indicated with a red/black edge line. The graph starts at the first layer, not at the embedding layer.
@@
DC (Dundee corpus) FPGD Kennedy et al. (2003) 16.49 17.75 17.41 15.13 10.44 17.10 16.32 15.39 13.53 10.49
NS (Natural Stories Corpus) SPR Futrell et al. (2021) 23.59 23.75 21.16 13.37 7.11 16.88 14.53 11.36 8.14 6.32
MAZE Boyce and Levy (2023) 1.63 4.26 8.22 18.19 31.81 9.70 17.56 24.15 32.86 39.63
ZuCO FPGD Hollenstein et al. (2018) 34.84 35.39 31.97 23.08 9.75 30.48 27.16 22.56 17.29 8.77
N400 Hollenstein et al. (2018) 0.06 0.10 0.15 0.20 0.15 0.20 0.32 0.34 0.29 0.18
UCL SPR Frank et al. (2013) 24.51 23.35 18.84 7.80 1.81 15.78 8.92 4.87 2.53 1.27
FPGD Frank et al. (2013) 22.62 26.24 25.12 15.55 5.02 16.28 14.48 11.87 9.47 5.57
N400 Frank et al. (2015) 57.45 33.30 14.01 12.89 32.26 11.31 6.12 16.19 29.49 37.11
Fillers in Vasishth et al. (2010) SPR Vasishth et al. (2010) 7.37 12.11 15.31 15.82 15.75 8.60 10.47 11.36 11.86 13.33
FPGD Vasishth et al. (2010) 8.93 8.07 8.44 8.98 12.01 8.94 10.91 12.91 13.81 14.00
MAZE Hahn et al. (2022) 4.72 2.99 7.53 35.86 83.16 9.96 28.27 52.00 73.38 88.64
Michaelov+, 2024 N400 Michaelov et al. (2024b) 1.07 1.79 2.37 2.10 1.13 0.95 1.51 1.70 1.38 0.99
Federmeier+, 2007 N400 Federmeier et al. (2007) 0.85 3.31 9.87 23.22 27.72 1.49 5.22 13.06 24.48 28.71
W&F, 2012 N400 Wlotko and Federmeier (2012) 0.44 0.24 0.09 0.08 0.13 0.51 0.27 0.12 0.05 0.11
Hubbard+, 2019 N400 Hubbard et al. (2019) 0.18 0.19 0.24 0.30 0.22 0.11 0.12 0.22 0.36 0.33
S&F,2022 N400 Szewczyk and Federmeier (2022) 0.13 0.18 0.39 1.02 1.36 0.16 0.33 0.77 1.29 1.42
Szewczyk+, 2022 N400 Szewczyk et al. (2022) 1.61 3.32 5.07 7.96 8.40 2.12 3.58 5.52 8.10 8.93
3.2 Language Models
We evaluated 21 open-source LMs including bilion-scale ones: GPT-2 (124M, 355M, 774M, and 1.5B parameters) Radford et al. (2019), OPT (125M, 1.3B, 2.7B, 6.7B, 13B, 30B, and 66B parameters) Zhang et al. (2022), and Pythia (14M, 31M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, and 12B parameters) Biderman et al. (2023). For tuned-lens experiments (§ 3.4), we used 14 of these models based on the availability of pre-trained tuned lenses. [^3] GPT-2 124M, 774M, 1.5B; OPT 125M, 1.3B, 6.7B; and Pythia 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B. The number of internal layers in our models ranges from 6 to 64. Results for surprisal from the embedding layer are excluded.
3.3 Psychometric Predictive Power
We evaluated the ability of surprisal values $\mathrm{surprisal}(w_{t})\in\mathbb{R}_{\geq 0}$ to predict word-by-word human reading data $\mathrm{Cost}(w_{t})\in\mathbb{R}$ using regression models with several baseline features:
$\mathrm{Cost}(w_{t})$ $\sim\mathrm{surprisal}(w_{t})+\mathrm{surprisal}(w_{t-1})$
$+\mathrm{surprisal}(w_{t-2})$
$+\mathrm{length}(w_{t})+\mathrm{freq}(w_{t})$
$+\mathrm{length}(w_{t-1})+\mathrm{freq}(w_{t-1})$
$+\mathrm{length}(w_{t-2})+\mathrm{freq}(w_{t-2})\;\;\mathrm{.}$ (1)
$\mathrm{surprisal}(w_{t})$ $=-\log p(w_{t}|\bm{w}_{<t})\;\;\mathrm{.}$ (2)
Here, $w_{t}$ is $t$-th word in a text, and $\bm{w}_{<t}=[w_{1},\cdots,w_{t-1}]$ is its context. Length and unigram frequency of words are included as baseline factors, [^4] Word length follows character number, and word frequency is estimated with word_freq Speer (2022). and $\mathrm{surprisal}(w_{t-1})\in\mathbb{R}_{\geq 0}$ and $\mathrm{surprisal}(w_{t-2})\in\mathbb{R}_{\geq 0}$ are also added to account for spillover effects. This model design follows recent studies Wilcox et al. (2023b); Pimentel et al. (2023), [^5] For N400 data, we added baseline amplitude feature.. We used statsmodels package Seabold and Perktold (2010).
We trained linear regression models both with and without $\mathrm{surprisal}(w_{t})$, and we report the difference in the goodness of fit measured by log-likelihood ($\Delta$ log-likelihood; $\Delta$LL). This indicates how strongly surprisal contributes to predicting human sentence processing data, i.e., higher $\Delta$LL is better. The key question is which LM layer computes surprisal with better $\Delta$LL.
Note that we used the same baseline factors across all datasets/models for a fair comparison, except for electrode random effects for Michaelov et al. (2024b)’s EEG data. As is common in preprocessing, we exclude data points with zero SPR/FPGD/MAZE. Human data for each token in the corpus were averaged across subjects prior to analysis, following recent practices Pimentel et al. (2023); Oh and Schuler (2023b); Kuribayashi et al. (2024); de Varda et al. (2024).
Figure 3: Scaling effect between model size (parameter counts in log scale) and $\Delta$LL. Each marker corresponds to each LM’s $\Delta$LL score from its best layer (red edge) or the last layer (black edge). The regression lines show the scaling effects, and the red line is for best-layer’s $\Delta$LL while the grey one is for the last layer’s. The maker type (shape/size/color) follows the legend in Figure 1. Pythia 14M, 31M, and 70M models are excluded as outliers in the X-axis, but including them does not alter the conclusion.
3.4 Probabilities from internal layers
Let us begin with the simplest method of logit-lens nostalgebraist (2020) to extract the probability of a word $w_{t}$ from model internals. Given a $d$-dimentaional internal representation $\bm{h}_{l,t}\in\mathbb{R}^{d}$ at the $l$-th layer and time step $t$, the probability of a word $w_{t}$ in its context $\bm{w}_{<t}$ is obtained as follows:
$=\mathrm{softmax}(\mathrm{LayerNorm}(\bm{h}_{l,t})\bm{W}_{U})[\mathrm{id}(w_{t})]\mathrm{,}$ (3)
where $\bm{W}_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is an unembedding matrix obtained from LM’s output layer, and $|\mathcal{V}|\in\mathbb{R}$ is model’s vocabulary size. Simply put, the internal representaion $\bm{h}_{l,t}$ is mapped into output vocabulary space by applying $\bm{W}_{U}$ (i.e., skipping subsequent layers [$\bm{h}_{l+1,t},\cdots,\bm{h}_{\mathrm{last},t}$]), and next-word probability is obtained in that space. The obtained probability (Eq. 3) is first transformed to surprisal (Eq. 2), and then used in the regression model to predict human reading data (Eq. 1). $\mathrm{LayerNorm}(\cdot)$ in Eq. 3 is the layer normalization at the last layer, and [$\mathrm{id}(w_{t})$] extracts the probability for $w_{t}$666 If a word is split into multiple subwords, accumulated surprisal is used. See Eq.2 in Kuribayashi et al. (2021). from the probability distribution over $\mathcal{V}$, obtained through the $\mathrm{softmax}(\cdot):\mathbb{R}^{|\mathcal{V}|}\to{[0,1]}^{|\mathcal{V}|}$ function.
Building on the logit-lens, Belrose et al. (2023) extended the method into tuned-lens to handle the potential representational drifts through layers. This technique introduces an additional linear transformation for each layer $l$ to mitigate the mismatch between the representation spaces of the $l$-th layer and the last layer:
$p(w|\bm{w}_{<t};\bm{h}_{l,t})$
$=\mathrm{LogitLens}(\bm{h}_{l,t}\bm{W}_{l}+\bm{b}_{l})[\mathrm{id}(w)]\mathrm{,}$ (4)
where $\bm{W}_{l}\in\mathbb{R}^{d\times d}$ and $\bm{b}_{l}\in\mathbb{R}^{d}$ are additionally trained to align the output of logit-lens with the last layer’s next-word probability distribution on additional LM pretraining data. [^7] We used publicly available tuned-lens: https://huggingface.co/spaces/AlignmentResearch/tuned-lens/tree/main. Their actual implementation used the representation of $\bm{h}_{l,t}\bm{W}^{\prime}_{l}+\bm{h}_{l,t}+\bm{b}_{l}=\bm{h}_{l,t}(\bm{W}^{\prime}_{l}+\bm{1})\bm{b}_{l}$, but we omit the identity matrix $\bm{1}\in\mathbb{R}^{d\times d}$ in Eq. 4 by overriding $\bm{W}_{l}=\bm{W}^{\prime}_{l}+\bm{1}$. Notably, we do not fine-tune any part of the LMs for human data; instead, we observe the emerging correlations between next-word probabilities and human reading measures.
4 Experimental results
Figures 2 (§ 4.1) and 3 (§ 4.2) summarize our main findings, with comprehensive results in Table 1.
4.1 Last layer does not yield the best $\Delta$LL
We first revisit the experimental settings (DC and NS datasets) of Oh and Schuler (2023b). The line graphs in Figure 2 depict the layer-wise $\Delta$LL for each LM. A consistent pattern emerges in the reading behavior (FPGD/SPR) modeling: the $\Delta$LL decreases toward the final layer (the rightmost, less-transparent, large markers), indicating that the last layer often yields the lowest score compared to the internal layers of the same model. These results challenge the assumption, widely adopted in existing studies, that the last layer is the most reliable indicator of an LM’s cognitive plausibility.
Table 1 presents a detailed breakdown of all the datasets, averaging $\Delta$LL scores across relative layer positions (e.g., 0–0.2 for the first 20% of layers). As shown in Figure 2, the best-performing layer is often not the final one (0.8–1.0). The optimal layer varies based on the type of measurement and stimuli. For instance, SPR and FPGD data are best modeled by earlier layers, whereas MAZE processing times and N400 signals are better captured by later layers (as motivated in § 2.3). Furthermore, the optimal layer for the UCL dataset is among the earlier layers, while for the Fillers dataset, it lies among the later layers (perhaps associated with the complexity of stimuli sentences). The layer–measure intearction are further analyzed in § 5.2. Notably, logit-lens and tuned-lens yilded geenrally consistent patterns.
We revisit the question with our extended focus on model internals: what kind of LMs yield the best $\Delta$LL from their internals? As the field is particularly interested in the relationship with model scaling Goodkind and Bicknell (2018); Oh and Schuler (2023b), we examine the relationship between LM parameter size (x-axis) and $\Delta$LL (y-axis) for two scenarios: (1) using the last layer’s $\Delta$LL (grey lines), reproducing previous findings, and (2) considering the best $\Delta$LL layer identified in this study (red lines).
@@ 4.2 Revisiting LM-scaling effects in cognitive modeling with internal layers
5 Analyses
Logit-lens (win rate) Tuned-lens (win rate)
GPT2 OPT OPT OPT OPT OPT OPT PT PT PT PT PT GPT2 OPT OPT PT PT PT PT
Data XL 1.3B 2.7B 6.7B 13B 30B 66B 1B 1.4B 2.8B 6.9B 12B XL 1.3B 6.7B 1.4B 2.8B 6.9B 12B
DC FPGD Kennedy et al. (2003) 0.80 0.80 0.82 0.76 0.73 0.76 0.78 0.00 0.32 0.36 0.73 0.73 0.73 0.80 0.67 0.64 0.58 0.55 0.54
NS SPR Futrell et al. (2021) 0.82 0.80 0.85 0.76 0.73 0.76 0.85 0.47 0.52 0.45 0.21 0.41 0.55 0.76 0.70 0.56 0.39 0.36 0.41
ZuCO FPGD Hollenstein et al. (2018) 0.80 0.84 0.88 0.79 0.73 0.78 0.86 0.76 0.72 0.70 0.85 0.84 0.80 0.84 0.76 0.72 0.67 0.61 0.65
UCL SPR Frank et al. (2013) 0.78 0.80 0.79 0.76 0.76 0.78 0.80 0.71 0.52 0.64 0.70 0.70 0.73 0.80 0.61 0.52 0.36 0.42 0.35
UCL FPGD Frank et al. (2013) 0.94 0.88 0.82 0.79 0.83 0.88 0.83 0.59 0.56 0.58 0.70 0.73 0.90 0.92 0.91 0.96 0.48 0.73 0.73
Table 2: How likely $\Delta$LL from internal layers outperformed the previous best $\Delta$LL (achieved within the same model family, relying on their last layers). The results are focused on billion-scale models and behavioral data with somewhat drastic flips in LM-scaling effects for $\Delta$LLs. “PT” denotes “Pythia.”
5.1 How easily can good layers be found?
To assess how many layers yield a good $\Delta$LL, we analyze the amount of internal layers that outperform the previously best $\Delta$LL achieved by the last layer within the same model family. Table 2 presents the win rate of internal layers’ $\Delta$LL against the respective previous best score. The win rate is typically around 80%, indicating a significant proportion of internal layers achieved good $\Delta$LL scores. These findings suggest that the cognitive plausibility of LLMs has been underestimated and that our argument (Figure 3) was not based on specific outlier layers but reflected a broader trend across many internal layers.
5.2 Layer depth and human measures
We observed systematic tendencies in the relationship between layer depth and human measurement methods. For instance, FPGD aligns better with earlier layers, whereas N400 aligns better with later layers, as summarized in Table 1. To statistically validate this relationship, we trained a linear regression model to explain $\Delta$LL scores from our 15,833 experimental settings $s\in\{\mathrm{dataset}\}\times\{\mathrm{model}\}\times\{\mathrm{layer}\}$, using the following formula:
$\Delta\mathrm{LL}(s)$ $\sim\mathrm{stimuli}(s)+\mathrm{model}(s)+\mathrm{lens}(s)$
$+$ $\mathrm{layer\_depth}(s)+\mathrm{measure}(s)$
$+$ $\mathrm{layer\_depth}(s)\times\mathrm{measure}(s),$ (5)
where $\mathrm{stimuli}$ represents the source stimuli of the data (“Stimuli” column in Table 1), $\mathrm{model}$ encodes the model name, $\mathrm{layer\_depth}$ is the depth of the layer where the $\Delta$LL is obtained, $\mathrm{measure}$ encodes the human measurement method (“Measure” column in Table 1), and $\mathrm{lens}$ indicates whether the logit-lens or tuned-lens was used. The term $\mathrm{layer\_depth}\times\mathrm{measure}$ captures the interaction between effective layer depth and human measures, which is of interest. Note that $\mathrm{measure}$ is a categorical variable, with $\mathrm{FPGD}$ serving as the dummy category.
[…]
It is worth noting that we also conducted an exploratory analysis with other human measures (e.g., P600, second-pass gaze duration) on the UCL corpus (Table 4 in Appendix). For such an extended scope, we cannot confirm the implied relationship between fast–slow human responses and early–late LM layers, and this appears to hold primarily for well-established human measures of SPR, FPGD, N400, and MAZE.
Logit-lens ($\Delta$LL) Tuned-lens ($\Delta$LL)
DC FPGD Kennedy et al. (2003) 1.49 1.71 1.76 1.63 1.20 1.59 1.60 1.65 1.58 1.30
NS SPR Futrell et al. (2021) 0.72 0.68 0.67 0.55 0.46 0.24 0.24 0.30 0.41 0.53
MAZE Boyce and Levy (2023) 0.45 0.46 0.55 0.85 1.25 0.67 0.70 0.76 1.02 1.30
ZuCO FPGD Hollenstein et al. (2018) 34.84 35.39 31.97 23.08 9.75 30.48 27.16 22.56 17.29 8.77
N400 Hollenstein et al. (2018) 0.06 0.10 0.15 0.20 0.15 0.20 0.32 0.34 0.29 0.18
UCL SPR Frank et al. (2013) 3.40 2.73 2.13 1.06 0.23 0.52 0.20 0.10 0.11 0.15
FPGD Frank et al. (2013) 7.96 9.21 9.07 6.61 2.86 2.46 2.29 2.22 2.04 1.21
N400 Frank et al. (2015) 3.77 1.96 2.64 8.07 16.31 1.34 2.60 6.30 9.52 13.26
Fillers SPR Vasishth et al. (2010) 0.37 0.40 0.89 2.89 8.74 0.55 1.19 1.31 1.68 2.94
FPGD Vasishth et al. (2010) 0.26 0.22 0.21 0.62 1.50 0.37 0.32 0.27 0.71 1.19
MAZE Hahn et al. (2022) 4.78 3.46 2.05 5.77 12.19 0.49 0.94 3.43 5.33 6.13
(a) All tokens.
(b) Clause-final tokens.
Figure 4: Relationship between $\Delta$LL and relative layer depth (lower is shallower) for each human measure. Different measures are associated with different layers; for example, good $\Delta$LLs for FPGD are achieved in earlier layers, while those for MAZE are in the latter layers.
5.3 Are results biased by targeted tokens?
A potential confound in § 5.2 stems from differences in targeted tokens for different human measures. For instance, N400 data are typically recorded only at sentence-final tokens to preprocess continuous, time-series EEG data, potentially introducing biases specific to sentence-final tokens, such as wrap-up effects Just and Carpenter (1980); Rayner et al. (2000); Meister et al. (2022). In contrast, behavioral measures are recorded across tokens within a sentence. To address this potential confound, we conducted additional experiments targeting sentence/clause-final tokens for all measures, even for SPR, FPGD, and MAZE. Sentence/clause-final tokens are ideitified using a constituency parser Kitaev et al. (2019); Kitaev and Klein (2018) and punctuations (e.g., “.”, “,”), following Meister et al. (2022).
[…]
5.4 Cross-lingual experiments
We also examined the generality of our findings across languages by following the experiments of de Varda and Marelli (2023), using FPGD data in 13 languages, recorded in MECO Siegelman et al. (2022) and five variants of multilingual XGLMs Lin et al. (2022). See detailed settings and results in Appendix A (Table 7 and Figure 7).
In summary, the best layer was typically not exactly the last layer (85%=55/66 of the settings), aligned with our main results. However, compared to our main results in Table 1, the best layers leaned toward later depths. In addition, LM-scaling effects are still negative in several langauges. These mixed results might be biased by the report that multilingual LMs process languages within the English subspace in middle layers Wendler et al. (2024), leading to reduced alignment with target languages. Future work will explore scaling experiments with monolingual LMs, as suitable models to test scaling effects across languages are currently unavailable.
6 When are earlier layers advantageous?
We lastly explore when and why earlier layers’ sueprisal aligns better with human reading data.
6.1 Residual error analysis

### v2

@@ 4.1 Human data
To minimize confounding factors between stimulus data and human measures, we included datasets with multi-layered annotations across multiple human measures. These include the Natural Stories Corpus Futrell et al. (2021) with SPR [^7] We use the version (2025-05-12) without the misalignment problem (see https://github.com/languageMIT/naturalstories). and MAZE data Boyce and Levy (2023), ZuCO corpus Hollenstein et al. (2018) with FPGD and N400 data, UCL Corpus Frank et al. (2013) annotated with SPR, FPGD, and N400 data Frank et al. (2015), and filler sentences from Hahn et al. (2022) annotated with SPR, FPGD, and MAZE data Vasishth et al. (2010); Hahn et al. (2022). In particular, the FPGD and N400 data in ZuCO were simultaneously recorded from the same human subjects, which likely minimized confounding factors. As is common in preprocessing, we exclude data points with zero SPR/FPGD/MAZE value. Human data for each token in the corpus were averaged across subjects prior to analysis, following recent practices Pimentel et al. (2023); Oh and Schuler (2023b); Kuribayashi et al. (2024); de Varda et al. (2024).
4.2 Language models
We evaluate 30 open-source LMs including bilion-scale ones: GPT-2 (124M, 355M, 774M, and 1.5B parameters) Radford et al. (2019), OPT (125M, 1.3B, 2.7B, 6.7B, 13B, 30B, and 66B parameters) Zhang et al. (2022), Pythia (14M, 31M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, and 12B parameters) Biderman et al. (2023), Qwen2.5 (0.5B, 1.5B, 3B, 7B, 14B, 32B, 72B), and Llama-3.1 (8B and 70B). See Appendix B for details. For tuned-lens experiments (Section˜3.1), we use 14 of these models based on the availability of pre-trained tuned lenses. [^8] We used parameters released in https://huggingface.co/spaces/AlignmentResearch/tuned-lens/tree/main. Specifically, we used GPT-2 124M, 774M, 1.5B; OPT 125M, 1.3B, 6.7B; and Pythia 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B, 12B, in the tuned-lens experiments. Notably, their actual implementation used the representation of $\bm{h}_{l,t}\bm{W}^{\prime}_{l}+\bm{h}_{l,t}+\bm{b}_{l}=\bm{h}_{l,t}(\bm{W}^{\prime}_{l}+\bm{1})+\bm{b}_{l}$, but we omit the identity matrix $\bm{1}\in\mathbb{R}^{d\times d}$ in Eq. 2 by overriding $\bm{W}_{l}=\bm{W}^{\prime}_{l}+\bm{1}$. The number of internal layers in our models ranges from 6 to 80. Results for surprisal from the embedding layer are excluded as these generally show a bad fit with human data.
(a) Logit-lens
(b) Tuned-lens
Figure 3: Scaling effect between model size (parameter counts in log scale) and $\Delta$LL. Each marker corresponds to each LM’s $\Delta$LL score from its best layer (red edge) or the last layer (black edge). The regression lines show the scaling effects, and the red line is for best-layer’s $\Delta$LL while the grey one is for the last layer’s. The maker type (shape/size/color) follows Figure 1.
5 Experimental results
Figures 2 (Section˜5.1) and 3 (Section˜5.2) summarize our main findings, with comprehensive results in Table 1.
5.1 Last layer does not yield the best $\Delta$LL
We first revisit the experimental settings (DC and NS datasets) of Oh and Schuler (2023b). The top two line graphs in Figure 2 depict the layer-wise $\Delta$LL for each LM. A consistent pattern emerges in FPGD and SPR data: the $\Delta$LL decreases toward the final layer (the rightmost, less-transparent, large markers), indicating that the last layer often yields the lowest score compared to the internal layers of the same model. These results challenge the assumption, widely adopted in existing studies, that the last layer is the most reliable indicator of an LM’s cognitive plausibility.
At the same time, changing the human measurement, i.e., from SPR to MAZE in the NS data (the second and third rows in Figure 2), flips the trend. This indicates that the aligned layers can easily change when using different human measurement methods, and suggests that different aspects of human sentence processing correlate with different LM layers, depending on the human measurement method. Such measurement–layer interaction effect is further looked into in Section˜6.2.
Table 1 presents a detailed breakdown of all the datasets, averaging $\Delta$LL scores across relative layer positions (e.g., 0–0.2 for the first 20% of layers). Table 1 includes the results with logit-lens and tuned-lens, which yielded generally consistent patterns. As shown in Figure 2, the best-performing layer is often not the final one (0.8–1.0), and the optimal layer varies based on the type of measurement and stimuli. For instance, SPR and FPGD data are best modeled by earlier layers, whereas MAZE processing times and N400 signals are better captured by later layers (as motivated in Section˜2.3). Furthermore, there are also stimulus-dependent biases — the optimal layer for the UCL dataset is among the earlier layers, while for the Fillers dataset, it lies among the later layers (perhaps associated with the specific complexity of the stimulus).
We revisit the question with our extended focus on model internals: what kind of LMs yield the best $\Delta$LL from their internals? As the field is particularly interested in the relationship with model scaling Goodkind and Bicknell (2018); Oh and Schuler (2023b), we examine the relationship between LM parameter size (x-axis) and $\Delta$LL (y-axis) for two scenarios: (1) using the last layer’s $\Delta$LL (grey lines), reproducing previous findings; and (2) considering the best $\Delta$LL layer identified in this study (red lines).
[…]
GPT2 OPT Pythia Qwen Llama-3
Data XL 1.3B 2.7B 6.7B 13B 30B 66B 1B 1.4B 2.8B 6.9B 12B 3B 7B 32B 72B 8B 70B
DC FPGD Kennedy et al. (2003) 0.80 0.80 0.82 0.76 0.73 0.76 0.78 0.00 0.32 0.36 0.73 0.81 0.10 0.26 0.72 0.94 0.89 0.73
NS SPR Futrell et al. (2021) 0.82 0.80 0.82 0.76 0.76 0.78 0.82 0.65 0.60 0.64 0.73 0.95 0.93 0.95 0.93 0.97 0.98 0.73
ZuCO FPGD Hollenstein et al. (2018) 0.80 0.84 0.88 0.79 0.73 0.78 0.86 0.65 0.60 0.55 0.76 0.97 0.97 0.97 0.98 0.97 0.98 0.78
UCL SPR Frank et al. (2013) 0.78 0.80 0.79 0.76 0.76 0.78 0.80 0.71 0.52 0.64 0.70 0.97 0.97 0.97 0.98 0.97 0.99 0.70
UCL FPGD Frank et al. (2013) 0.94 0.88 0.82 0.79 0.83 0.88 0.83 0.59 0.56 0.58 0.70 0.97 0.93 0.97 0.94 0.97 0.99 0.73
Table 2: How likely $\Delta$LL from internal layers outperformed the previous best $\Delta$LL (achieved within the same model family, relying on their last layers). The results are focused on billion-scale models and behavioral data with somewhat drastic flips in LM-scaling effects for $\Delta$LLs.
6 Analyses
We conduct several follow-up analyses to support and clarify the findings reported in Section˜5.
6.1 How easily can good layers be found?
[…]
6.2 Layer depth and human measures
We observed systematic tendencies in the relationship between layer depth and human measurement methods. For instance, FPGD aligns better with earlier layers, whereas N400 aligns better with later layers, as summarized in Table 1. To statistically validate this relationship, we trained a linear regression model to explain $\Delta$LL scores from our 23,154 experimental settings $\{\mathrm{dataset}\}\times\{\mathrm{model}\}\times\{\mathrm{layer}\}$ with the following features for each setting $s$: $\mathrm{stimuli}(s)$, $\mathrm{model}(s)$, $\mathrm{lens}(s)$, $\mathrm{layer\_depth}(s)$, $\mathrm{measure}(s)$, $\mathrm{layer\_depth}(s)$:$\mathrm{measure}(s)$}. Here, $\mathrm{stimuli}$ represents the source stimuli of the data (“Stimuli” column in Table 1), $\mathrm{model}$ encodes the model name, $\mathrm{layer\_depth}$ is the depth of the layer where the $\Delta$LL is obtained, $\mathrm{measure}$ encodes the human measurement method (“Measure” column in Table 1), and $\mathrm{lens}$ indicates whether the logit-lens or tuned-lens was used. The term $\mathrm{layer\_depth}\times\mathrm{measure}$ captures the interaction between effective layer depth and human measures, which is of interest. Note that $\mathrm{measure}$ is a categorical variable, with $\mathrm{SPR}$ serving as the dummy category.
[…]
(a) All tokens.
(b) Clause-final tokens.
Figure 4: Relationship between $\Delta$LL and relative layer depth for each human measure. Different measures are associated with different layers; for example, good $\Delta$LLs for FPGD are achieved in earlier layers, while those for MAZE are in the latter layers.
@@ 6.2 Layer depth and human measures
DC FPGD Kennedy et al. (2003) 1.24 1.37 1.42 1.31 1.00 1.59 1.60 1.65 1.58 1.30
NS SPR Futrell et al. (2021) 0.45 0.38 0.33 0.29 0.21 0.16 0.12 0.12 0.18 0.25
MAZE Boyce and Levy (2023) 0.35 0.36 0.42 0.62 0.84 0.67 0.70 0.76 1.02 1.30
ZuCO FPGD Hollenstein et al. (2018) 38.10 38.13 35.59 29.82 15.94 30.48 27.16 22.56 17.29 8.77
N400 Hollenstein et al. (2018) 0.07 0.12 0.15 0.18 0.16 0.20 0.32 0.34 0.29 0.18
UCL SPR Frank et al. (2013) 3.09 2.53 2.12 1.21 0.50 1.76 0.89 0.50 0.27 0.15
FPGD Frank et al. (2013) 7.51 7.89 8.33 6.67 3.56 4.85 4.86 4.60 4.26 2.66
N400 Frank et al. (2015) 3.58 1.66 1.99 5.20 11.95 1.12 3.85 8.31 12.75 18.17
Fillers SPR Vasishth et al. (2010) 0.28 0.33 0.62 1.86 5.78 0.88 2.19 3.32 5.83 10.31
FPGD Vasishth et al. (2010) 0.30 0.25 0.26 0.59 1.29 0.28 0.42 0.52 1.40 2.04
MAZE Hahn et al. (2022) 3.54 2.41 1.51 3.83 8.99 1.15 1.94 5.52 10.64 12.53
6.3 Are results biased by targeted tokens?
A potential confound in Section˜6.2 stems from differences in targeted tokens for different human measures. For instance, N400 data are typically recorded only at sentence-final tokens to preprocess continuous, time-series EEG data, potentially introducing biases specific to sentence-final tokens, such as wrap-up effects Just and Carpenter (1980); Rayner et al. (2000); Meister et al. (2022). In contrast, behavioral measures are recorded across tokens within a sentence. To address this potential confound, we conducted additional experiments targeting sentence/clause-final tokens for all measures, even for SPR, FPGD, and MAZE. Sentence/clause-final tokens are ideitified using a constituency parser Kitaev et al. (2019); Kitaev and Klein (2018) and punctuations (e.g., “.”, “,”), following Meister et al. (2022).
@@ 6.3 Are results biased by targeted tokens?
6.4 When are earlier layers advantageous?